# PerfBench: Can Agents Resolve Real-World Performance Bugs?

### v2

@@ IV-C Agent Configurations
Each agent is allowed a maximum of 100 steps per task to mimic realistic usage without being too expensive and time-consuming, while still providing ample opportunity for the agent to iterate over the problem. Tasks are executed in isolated Docker containers to ensure reproducibility and prevent interference between runs. We share a sample code for a docker image in the Appendix.
IV-D A Perf Agent
@@
IV-D1 Performance-Aware & Benchmarking Instructions
@@ 6. Finally explain the fix you implemented and output
Figure 4: Prompt template differences between baseline OpenHands (light red) and OpenHands-Perf-Agent (light green). The performance-aware version replaces error reproduction with benchmark creation and emphasizes the high-level sequence of steps we expect the agent to follow to fix a performance bug.
  • Success Case: When benchmarks execute successfully, i.e. we see a table at the end of the output, we extract only the summary table (as shown in Figure 3) containing key performance metrics and discard verbose diagnostic information preceding it.
  • Error Case: When benchmarks fail, we preserve the complete output to enable the agent to diagnose and rectify issues.
@@
V Benchmark Results
V-A Overall Performance
Table III presents the main experimental results of our two agent configurations on PerfBench with the LLM models we use.
TABLE III: Overall performance of agents on PerfBench
@@ V-A Overall Performance
TABLE IV: Resolution rates by high-level performance bug categories in Table II of our Baseline OpenHands and OpenHands-Perf-Agent configurations with Claude Sonnet 4
@@
The results reveal significant challenges in performance bug fixing for current LM agents. The baseline OpenHands agent achieves <4% success rate across both models, substantially lower than typical performance on functional bug fixing benchmarks such as SWE-bench Verified (>60% [8]). This dramatic performance gap highlights the fundamental differences between functional and performance bug fixing tasks.
@@
V-B Performance Metrics Analysis
To provide deeper insights into the nature of performance improvements achieved by our agents, we analyze the change in CPU and memory metrics captured by our harness for successful fixes. Figure 5 presents the distribution of absolute and relative improvements in memory allocation and CPU utilization compared to the original code.
Memory allocation reductions (Figure 5 demonstrate the more prominent improvements, with all successful fixes typically having more than 20% improvement to allocations, with several having improvements on the order of KBs and MBs. Unlike Memory, CPU utilization improvements (Figure 5 are typically in the range from 0 to 40%, but are on the order of microseconds or milliseconds. The highest CPU improvements correspond to fixes addressing algorithmic inefficiencies. An example for one such algorithmic change suggested by agent is provided in the Appendix (Figure 8).
@@ V-B Performance Metrics Analysis
V-C Performance by Bug Category
Table IV breaks down the performances of OpenHands-Perf-Agent and baseline OpenHands (with Claude Sonnet 4) across the different performance bug categories identified in the high-level categories we found in Section III-E.
@@ V-C Performance by Bug Category
Notably, the baseline agent fails completely on most of the categories, achieving only a handful of successful runs in memory management and algorithmic issues. This suggests that without explicit performance-aware instruction, agents struggle to identify and address more complex performance patterns beyond the most obvious inefficiencies. Even the OpenHands-Perf-Agent configuration struggles with most categories of issues such as Build & Analysis Tools, Concurrency, etc., showing that there is room for much improvement.
VI Limitations
While PerfBench provides valuable insights into agent capabilities for performance optimization, our work has several limitations that should be considered when interpreting results and directions for future work.
Performance vs. Correctness Trade-off: Our evaluation harness primarily focuses on measurable performance improvements rather than assessing whether agents implement the correct fix for the underlying performance issue discussed in the GitHub issue. An agent may achieve performance gains through alternative optimizations that differ from the developer’s intended solution, yet still receive full credit in our evaluation. While such improvements may be practically valuable, they may not demonstrate true understanding of the root cause identified by the original developer. Addressing this limitation would require sophisticated semantic analysis, potentially using LLM-as-a-Judge critics to evaluate solution correctness, which we leave for future exploration.
Dependency on Benchmarking: Our evaluation relies entirely on agent-generated benchmarks, creating a potential weakness where poorly designed benchmarks may fail to capture the true performance characteristics of the issue. Agents that generate inadequate benchmarks may receive false negatives even with correct fixes. However, we would like to note that writing benchmarks in .NET is a relatively easy task for today’s LLMs and easier still compared to writing unit tests which require deeper understanding of the code and coming up with test cases that exercise crucial code paths. On the other hand benchmarks are simply measurement tools that don’t require such deep analysis. Our evaluation framework requires agents to use the BenchmarkDotNet framework for performance measurement. While this reflects standard .NET practices, it constrains agents to a specific benchmarking methodology and may disadvantage approaches that rely on alternative but correct performance measurement techniques or custom profiling solutions.
@@ VI Limitations
Focus on CPU and Memory: Our evaluation primarily considers improvements in execution time and memory allocation without capturing other important performance dimensions such as network usage, disk I/O usage, or end user-perceived latency. This narrow focus may miss optimization and bug-fix opportunities that improve the overall system performance, but do not manifest as improvements in our benchmarking.

### v3

@@ 4.3. Agent Configurations
Each agent is allowed a maximum of 100 steps per task to mimic realistic usage without being too expensive and time-consuming, while still providing ample opportunity for the agent to iterate over the problem. Tasks are executed in isolated Docker containers to ensure reproducibility and prevent interference between runs.
4.4. A Perf Agent
@@
4.4.1. Performance-Aware & Benchmarking Instructions
@@ 6. Finally explain the fix you implemented and output
Figure 4. Prompt template differences between baseline OpenHands (light red) and OpenHands-Perf-Agent (light green). The performance-aware version replaces error reproduction with benchmark creation and emphasizes the high-level sequence of steps we expect the agent to follow to fix a performance bug.
  • Success Case: When benchmarks execute successfully, i.e. we see a table at the end of the output, we extract only the summary table (shown in Figure 3) containing key perf metrics and discard verbose diagnostics preceding it.
  • Error Case: When execution fails, we preserve the complete output to enable agent to diagnose and rectify issues.
@@
5. Benchmark Results
5.1. Overall Performance
Table 4 presents the main experimental results of our two agent configurations on PerfBench with the LLM models we use.
Table 4. Overall performances of different agents and model configurations on PerfBench
@@ 5.1. Overall Performance
Table 5. Resolution rates by high-level performance bug categories in Table 3 of our Baseline OpenHands and OpenHands-Perf-Agent configurations with Claude Sonnet 4
@@
The results reveal significant challenges in performance bug fixing for current LM agents. The baseline OpenHands agent achieves <4% success rate across both models, substantially lower than typical performance on functional bug fixing benchmarks such as SWE-bench Verified (>60% (Wang et al., 2024)). This dramatic performance gap highlights the fundamental differences between functional and performance bug fixing tasks.
@@
5.2. Performance Metrics Analysis
To provide deeper insights into the nature of performance improvements achieved by our agents, we analyze the change in CPU and memory metrics for successful fixes. Figure 5 presents the distribution of absolute and relative improvements in memory allocation and CPU utilization compared to the original code.
Memory allocation reductions (Figure 5 demonstrate the more prominent improvements, with all successful fixes typically having more than 20% improvement to allocations, with several having improvements on the order of KBs and MBs. Unlike Memory, CPU utilization improvements (Figure 5 are typically in the range from 0 to 40%, but are on the order of microseconds or milliseconds. The highest CPU improvements correspond to fixes addressing algorithmic inefficiencies.
@@ 5.2. Performance Metrics Analysis
5.3. Performance by Bug Category
Table 5 breaks down the performances of OpenHands-Perf-Agent and baseline OpenHands (with Claude Sonnet 4) across the different performance bug categories identified in the high-level categories we found in Section 3.5.
@@ 5.3. Performance by Bug Category
Notably, the baseline agent fails completely on most of the categories, achieving only a handful of successful runs in memory management and algorithmic issues. This suggests that without explicit instructions, agents struggle to identify and address complex performance patterns beyond the most obvious inefficiencies. Even the OpenHands-Perf-Agent configuration struggles with other categories such as Build & Analysis Tools, Concurrency, etc., showing that there is room for much improvement.
6. Limitations
While PerfBench provides valuable insights into agent capabilities for perf optimization, our work has various limitations that must be considered.
Performance vs. Correctness Trade-off: Our evaluation harness primarily focuses on measurable performance improvements rather than assessing whether agents implement the correct fix for the underlying performance issue discussed in the GitHub issue. An agent may achieve performance gains through alternative optimizations that differ from the developer’s intended solution, yet still receive full credit in our evaluation. While such improvements may be practically valuable, they may not demonstrate true understanding of the root cause identified by the original developer.
Dependency on Benchmarking: Our evaluation relies entirely on agent-generated benchmarks, creating a potential weakness where poorly designed benchmarks may fail to capture the true performance characteristics of the issue. Agents that generate inadequate benchmarks may receive false negatives even with correct fixes. However, we would like to note that writing benchmarks in .NET is a relatively easy task for today’s LLMs and easier still compared to writing unit tests which require deeper understanding of the code and coming up with test cases that exercise crucial code paths. On the other hand benchmarks are simply measurement tools that don’t require such deep analysis.
@@ 6. Limitations
Focus on CPU and Memory: Our evaluation primarily considers improvements in execution time and memory allocation without capturing other important performance dimensions such as network usage, disk I/O usage, or end user-perceived latency.